# The Dance of the Sheared Eigenfunctions

### v1

@@ 4 Conclusions and final remarks
In this work, we considered two families of sheared potential wells: a family of linear monomials and a family of harmonic oscillators. In both cases, we have chosen sheared potentials that retained the same functional form as the initial potential for $x<0$ and $x\geq 0$, but with different coefficients. For both families, we computed not only the spectra but also the corresponding eigenfunctions. In order to compute these spectra it was necessary to solve transcendental equations with the aid of numerical methods. According to Felix Klein (1849-1925), “it is well known that the central problem of the whole of modern mathematics is the study of transcendental functions defined by differential equations[22]. Although this modern is something old, the pursuit of this kind of treatment, combined with the tools of special functions and modern computing software, is a reviving topic in physics and mathematics on the last few decades [23].
As expected, we verified that the energy spectrum did not remain constant during the shearing process. We showed that this is due to the variation of the eigenfunctions. The main message of our work is that analysing the sheared eigenfunctions is a key element to understand why the spectra change. Combining the eigenfunctions with considerations about the work performed we could understand why the spectra display a decrease in energy with increasing $\nu$ in the vicinity of $\nu=0.5$, and also why the variations in this region is more significant then around $\nu=1$. This helps us understand the close similarities displayed by the results for the linear and the harmonic wells and we believe that these features would also be present for other power-law potentials. A perspective of this paper is to deepen this connection, evaluating quantitatively the work done in the process.
Once this calculation is made, one could investigate what would be the energetic costs during other processes of changing the potential wells in order to find the one which would require a minimum work to link the initial and final potentials. This kind of investigation could be useful to establish a connection with the promising and effervescent area of shortcut to adiabaticity.
@@ Acknowledgments
All authors thank M. Asorey for insightful discussions. We also thank the Brazilian agencies CAPES, CNPq and FAPERJ for financial support. J.O.-C. is grateful to CAPES (Grant No. 88887.958122/2024-00). C.F. acknowledges funding from CNPq (Grants No. 308641/2022-1 and 408735/2023-6) and FAPERJ (Grant No. 204.376/2024). F.S.S.R. is grateful to CNPq (Grant No. 315730/2021-8).

### v2

@@ 4 Conclusions and final remarks
In this work, we considered two families of sheared potential wells: a family of linear monomials and a family of harmonic oscillators. In both cases, we have chosen sheared potentials that retained the same functional form as the initial potential for $x<0$ and $x\geq 0$, but with different coefficients. For both families, we computed not only the spectra but also the corresponding eigenfunctions. In order to compute these spectra it was necessary to solve transcendental equations with the aid of numerical methods. According to Felix Klein (1849-1925), “it is well known that the central problem of the whole of modern mathematics is the study of transcendental functions defined by differential equations[klein1893]. Although this modern is now somewhat old, the pursuit of this kind of treatment, combined with the tools of special functions and modern computing software, is a reviving topic in physics and mathematics on the last few decades [andrews1999special]. For the families of sheared potentials considered in this work, we showed that the energy spectrum in both families did not remain constant during the shearing process.
We showed that this is due to the variation of the eigenfunctions. The take-home message of our work is that analysing the sheared eigenfunctions is a key element to understand the change in the spectra. Combining the eigenfunctions with considerations about the work performed we could understand why the spectra display a decrease in energy with increasing $\nu$ in the vicinity of $\nu=0.5$, and also why the variation in this region is more significant then around $\nu=1$. This helps us understand the close similarities displayed by the results for the linear and the harmonic wells and we believe that these features would also be present for other power-law potentials. A perspective of this paper is to deepen this connection, evaluating quantitatively the work done in the process. Once this calculation is made, one could investigate what would be the energetic costs during other processes of changing the potential wells in order to find the one which would require a minimum work to link the initial and final potentials.
To investigate how the evolution of quantum systems can be affected by tunable external agents, whose influence are encoded in continuous parameters present in the corresponding hamiltonians, may provide useful mechanisms of controlling the time evolution of such systems The cases considered in this paper are particularly appealing for modelling asymmetric quantum wells, which are used, for instance, on the study of the quantum-confined Stark-effect [Miller1984] or for the Wannier-Stark localization in optical latices [Emin1987], in the case of the first monomial well. For Stark-related effects, the shearing parameter, which controls the inclination of the curve, is directly proportional to the intensity of the applied electric field. For the harmonic oscillator case, an interesting example is to consider a time-dependent frequency and analyse, with the aid of a parameter that characterizes how abrupt is the change in the frequency, the time evolution of the squeezing parameter for frequencies with different time dependences. One can even investigate a continuous route from an adiabatic change to a completely abrupt one [MartnezTibaduiza2021]. This kind of study may be useful in a rapidly developing area of physics called shortcut to adiabaticity [delCampo2013] and their relation with harmonic traps [Chen2010, Hardel2024]. In this sense, the present work provides some kind of control of quantum spectra and wavefunctions which deserves a deeper investigation, since the shearing process may be useful in a variety of situations.
@@ Acknowledgments
All authors thank M. Asorey for insightful discussions. We also thank the Brazilian agencies CAPES, CNPq and FAPERJ for financial support. J.O.-C. is grateful to FAPERJ (Master’s scholarship No. 201.879/2025). C.F. acknowledges funding from CNPq (Grants No. 308641/2022-1 and 408735/2023-6) and FAPERJ (Grant No. 204.376/2024). F.S.S.R. is grateful to CNPq (Grant No. 315730/2021-8).
Appendix A Résumé of some special functions
In this paper, we made use of two important special functions, namely the Airy and the $D$-parabolic cylinder functions. In this appendix, we discuss the principal features of these functions.
A.1 Airy functions
The Airy functions originate from the Airy (or Stokes) equation,
$\dfrac{d^{2}f}{dz^{2}}(z)-zf(z)=0\,.$ (28)
The general solution of this problem is a sum of the functions called Airy functions, named after the astronomer G. B. Airy, defined by
$\mathrm{Ai}(z)$ $=\dfrac{1}{\pi}\int_{0}^{\infty}\cos\left(\dfrac{t^{3}}{3}+zt\right)\,dt\,$, (29)
$\mathrm{Bi}(z)$ $=\dfrac{1}{\pi}\int_{0}^{\infty}\left\{\exp\left(-\dfrac{t^{3}}{3}+zt\right)+\sin\left(\dfrac{t^{3}}{3}+zt\right)\right\}\,dt\,.$ (30)
Note that as $\mathrm{Bi}$ diverges for $z\to\infty$, its contribution was discarded in Sec. 2. For analytical calculations, it is useful to write these functions in terms of Bessel functions (consider $x\geq 0$):
$\mathrm{Ai}(x)$ $=\frac{1}{\pi}\sqrt{\frac{x}{3}}K_{1/3}\left(\frac{2}{3}x^{3/2}\right)$ (31)
$\mathrm{Ai}(-x)$ $=\sqrt{\frac{x}{9}}\left[J_{1/3}\left(\frac{2}{3}x^{3/2}\right)+J_{-1/3}\left(\frac{2}{3}x^{3/2}\right)\right]$ (32)
$\mathrm{Bi}(x)$ $=\sqrt{\frac{x}{3}}\left[I_{1/3}\left(\frac{2}{3}x^{3/2}\right)+I_{-1/3}\left(\frac{2}{3}x^{3/2}\right)\right]$ (33)
$\mathrm{Bi}(-x)$ $=\sqrt{\frac{x}{3}}\left[J_{-1/3}\left(\frac{2}{3}x^{3/2}\right)-J_{1/3}\left(\frac{2}{3}x^{3/2}\right)\right]$ (34)
Expansions and power laws for the Airy functions can be found for example in [abramowitz]. The general behaviour of these functions is illustrated in Fig. (7)
Figure 7: General behaviour of the Airy functions and its derivatives. Note that the panel for the Bi function is in log-log scale.
A.2 D-Parabolic Cylinder functions
The parabolic cylinder functions originate from the differential equation
$\dfrac{d^{2}f}{dz^{2}}(z)+(az^{2}+bz+c)f(z)=0\,$, (35)
which surges naturally from Laplace’s equation when we separate its variables in parabolic cylindrical coordinates.
A particular case of this equation is the so-called Weber equation,
$\dfrac{d^{2}f}{dz^{2}}(z)+\left(a-\dfrac{1}{4}z^{2}\right)f(z)=0\,.$ (36)
The independent solutions of this equation are commonly expressed as [abramowitz]
$y_{1}(a;z)$ $=\exp(-z^{2}/4)\,{}_{1}F_{1}\left(\tfrac{1}{2}a+\tfrac{1}{4};\tfrac{1}{2};\tfrac{z^{2}}{2}\right)\,$, (37)
$y_{2}(a;z)$ $=z\exp(-z^{2}/4)\,{}_{1}F_{1}\left(\tfrac{1}{2}a+\tfrac{3}{4};\tfrac{3}{2};\tfrac{z^{2}}{2}\right)\,$, (38)
where $_{1}F_{1}(a;b;z)$ is the confluent hypergeometric function.
The parabolic cylinder functions are particular combinations of the previous equations, defined as
$U(a,z)$ $=\frac{1}{2^{x}\sqrt{\pi}}\left[\cos(x\pi)\Gamma\left(\tfrac{1}{2}-x\right)y_{1}(a,z)-\sqrt{2}\sin(x\pi)\Gamma(1-x)\,y_{2}(a,z)\right]\,$, (39)
$V(a,z)$ $=\frac{1}{2^{x}\sqrt{\pi}\Gamma(1/2-a)}\left[\sin(x\pi)\Gamma\left(\tfrac{1}{2}-x\right)y_{1}(a,z)+\sqrt{2}\cos(x\pi)\Gamma(1-x)\,y_{2}(a,z)\right]$ (40)
where $x=\tfrac{1}{2}a+\tfrac{1}{4}$.
In this paper, we are concerned about the particular case of the Weber equation
$\dfrac{d^{2}\psi}{dz^{2}}(z)+\left(\sigma+\dfrac{1}{2}-\dfrac{1}{4}z^{2}\right)\psi(z)=0\,$, (41)
which is Schrödinger’s equation for the harmonic oscillator when we identify $x=z\sqrt{\hbar/2m\omega}$ and $\sigma=\dfrac{\mathcal{E}}{\hbar\omega}-\dfrac{1}{2}$. In this case, following Whittaker [whittaker1990], the solutions are given in terms of the $D$-parabolic cylinder functions as given in Eq. 21, such that
$U(a,z)$ $=D_{-a-\frac{1}{2}}(z)\,$, (42)
$V(a,z)$ $=\frac{\Gamma\left(\tfrac{1}{2}+a\right)}{\pi}\left[\sin(\pi a)D_{-a-\frac{1}{2}}(z)+D_{-a-\frac{1}{2}}(-z)\right]\,.$ (43)
The advantage to write the $D$-parabolic cylinder function in terms of $U(a,z)$ is that it is relatively easy to expand $U(a,z)$ in a power series of $z$,
$U(a,z)=\frac{\sqrt{\pi}\,2^{-\frac{a}{2}-\frac{1}{4}}}{\Gamma\left(\frac{a}{2}+\frac{3}{4}\right)}-\frac{\sqrt{\pi}\,2^{-\frac{a}{2}+\frac{1}{4}}}{\Gamma\left(\frac{a}{2}+\frac{1}{4}\right)}z+\frac{\sqrt{\pi}\,2^{-\frac{a}{2}-\frac{5}{4}}}{\Gamma\left(\frac{a}{2}+\frac{3}{4}\right)}z^{2}-\cdots\,$, (44)
such that
$U(a,0)=\dfrac{\sqrt{\pi}2^{-a/2-1/4}}{\Gamma\left(\dfrac{a}{2}+\dfrac{3}{4}\right)}\quad\text{and}\quad U^{\prime}(a,0)=-\dfrac{\sqrt{\pi}2^{-a/2+1/4}}{\Gamma\left(\dfrac{a}{2}+\dfrac{1}{4}\right)}\,.$ (45)
In this way,
$D_{\frac{E}{\nu}-\frac{1}{2}}(0)=\dfrac{\sqrt{\pi}2^{E/2\nu-1/4}}{\Gamma\left(-\dfrac{E}{2\nu}+\dfrac{3}{4}\right)}\quad\text{and}\quad D^{\prime}_{\frac{E}{\nu}-\frac{1}{2}}(0)=-\dfrac{\sqrt{\pi}2^{E/2\nu+1/4}}{\Gamma\left(-\dfrac{E}{2\nu}+\dfrac{1}{4}\right)}\,.$ (46)
Taking $\nu=1$ we get the expressions used in Sec. 3.
Finally, a useful relation is that, when $\sigma\in\mathds{N}$, $D_{\sigma}(z)$ becomes proportional to a Hermite polynomial [bateman2006higher]:
$D_{\sigma}(z)=\mathrm{e}^{-z^{2}/4}2^{-\sigma/2}H_{\sigma}(z/\sqrt{2})\,.$ (47)
As the not sheared harmonic oscillator is such a case with $\sigma\in\mathds{N}$, the route of solving the problem with $D$-parabolic cylinder functions is almost never taken. In fact, almost all textbooks do not even mention this function (Merzbacher [merzbacher1998quantum] uses it but just in the context of a double harmonic oscillator).
Expansions and power laws for the $D$-parabolic cylinder function can be found for example in [KMAbadir1993]. The general behaviour of $D_{\sigma}(x)$ is illustrated in Fig. 8.
Figure 8: General behaviour of the $D$-parabolic function for real and imaginary arguments. For real arguments, the function is purely real and falls out rapidly (left-superior panel). For imaginary arguments, the imaginary part converges only if $\sigma$ is even (bottom panels), and the real part converges only if $\sigma$ is odd (right-superior panel).